# Structure-reinforced Transformer for Dynamic Graph Representation Learning with Edge Temporal States

### v2

@@ IV-C Experiment Results and Analyses
Following the experimental framework outlined in [10], we use dynamic link prediction [37] to evaluate the efficacy of the proposed RSGT model. Given a dynamic graph $\mathcal{G}$, we divide the snapshots from various time slices into training and test datasets. The training dataset $\mathcal{D}^{tr}$ encompasses all snapshots captured prior to time $tr$. These snapshots are utilized for training the representation learning models. Conversely, the snapshots captured post time $tr$ are allocated to the test dataset $\mathcal{D}^{te}$. Taking into consideration both the size and complexity of the datasets, we designate the dimension of node representation $d$ as 32 for the twi-Tennis, CollegeMsg, and sx-MathOF datasets, and 16 for the cit-HepTh dataset. During the testing phase, we initially generate temporal node representations using the trained model. These representations are subsequently input into a logistic regression classifier implemented via scikit-learn [^1] https://scikit-learn.org/ to predict the likelihood of a candidate edge being either positive or negative, assessed one snapshot at a time. For each test snapshot in $\mathcal{D}^{te}$, we adopt a 1:1 ratio of negative sampling to construct the complete test edge set. Here, 80% of the total test edges are utilized to train the classifier, while the remaining 20% are reserved for testing purposes. The scale factor $\alpha$ and the power factor $\beta$ are kept constant at 1 in all datasets. We conduct the experiment five times and report the mean classification accuracy and F1 scores, along with their respective standard deviations.
Comparison with competing baselines
@@
The experimental results for dynamic link prediction, as outlined in Table IV and Table IV, reveal a comprehensive comparison amongst the various baseline methods. Sub-optimal performance is indicated by darker shading, while the most striking results are highlighted in bold. Our proposed method, RSGT, demonstrates a consistently superior performance across all four experimental datasets when compared to the baseline techniques. This superiority is evident from the relative increments in Accuracy, which range from 8.50% to 16.45%. Similarly, the F1 scores also show considerable improvement, ranging from 7.44% to 14.6%. This assertion of RSGT’s exceptional effectiveness is thereby clearly substantiated by the empirical evidence.
Figure 3: Performance of different ablated variations of RSGT.
Analyzing from the viewpoint of various method categories, it is noticeable that dynamic graph approaches tend to outperform those that designed for static graphs. This suggests that the integration of temporal information is pivotal for comprehending the dynamic evolution of graphs, and consequently, aids the model in learning enhanced node representations. However, it is noteworthy that some static graph approaches including Node2vec and GraphSAGE, supersede dynamic graph approaches such as EvolveGCN, CoEvoSAGE, and CTDNE on the cit-HepTh dataset. This outcome arises primarily due to the drastic alterations in edges between consecutive snapshots within this dataset. A significant fraction of edges exist only within a singular time slice, posing a challenge for certain dynamic graph approaches to capture such extreme fluctuations.
Furthermore, when evaluating the baseline methods for dynamic graphs, we find that approaches designed for continuous dynamic graphs typically surpass those intended for discrete dynamic graphs. This superiority can be predominantly attributed to the potential loss of fine-grained temporal information during the discretization of graphs. As a result, discrete dynamic graph methods may perceive lesser information throughout the training phase, thereby imposing a limit on their performance capabilities.
TREND outperforms other baselines significantly owing to its design catering to continuous dynamic graphs, inherently preventing the loss of evolutionary information due to graph discreteness. Additionally, its GNN architecture makes it inductive, and the integration of the Hawkes process allows it to capture interesting effects between temporal edges. TREND also incorporates both edge and node dynamics, thus capturing the individual and collective characteristics of temporal edges and enabling a more accurate modeling of the temporal process.
Contrarily, RSGT showcases stable performance across all datasets, regardless of their scale and sparsity, and maintains high performance under varying train/test ratios, thus demonstrating its robustness and superior reductive learning capability. The optimal performance of RSGT on different datasets can be principally credited to two factors. First, node representation learning fundamentally involves a biased aggregation of connective nodes’ features along the edges. The temporal state of these edges directly influences the correlation between node pairs. Thus, explicitly modeling the edge temporal states can provide additional heuristic information via the weighted multi-relation dynamic graph, thereby enhancing the learning of node representations. Furthermore, RSGT can autonomously identify and mitigate the influence of erratic neighbors during the recurrent updates of node representations, resulting in consistently high performance across all datasets, regardless of their evolutionary patterns. Second, the proposed structure-reinforced graph transformer can capture the global semantic correlation between node pairs and extract the deep latent feature of graph topology and edge temporal states, leading to the generation of high-quality node representations.
Figure 4: Performance impact of parameters at different datasets.
IV-D Ablation Study
To evaluate the individual contributions of each component within the RSGT, we performed a performance comparison of dynamic link prediction across six ablated RSGT variants:
  1. 1.
    RSGT-T: This variant excludes varying edge types and does not reconstruct vanished edges.
  2. 2.
    RSGT-W: This variant assigns a fixed weight of 1 to different edge types, hence ignoring the long-term dynamics.
  3. 3.
@@ IV-D Ablation Study
    RSGT-TW: In this variant, the original dynamic graph is used directly without any explicit modeling of the temporal states of the edges.
  4. 4.
@@
With respect to learning edge temporal states, the findings from Figure 3a - Figure 3d suggest that the performance of RSGT-T and RSGT-W is roughly equivalent on the twi-Tennis, ColledgeMsg, and sx-MathOF datasets. This indicates that the weights or types of edges provide some of the temporal state information to the representation learning model, hence the inferior performance of RSGT-TW is expected. Interestingly, RSGT and RSGT-W demonstrate clear performance superiority over RSGT-T and RSGT-TW on the cit-HepTh dataset. This can be mainly attributed to the fact that most edges typically exist only for a single time slice, thereby making the long-term dynamics reflected in the edge weights have a minimal impact on the performance of RSGT-W for dynamic graph representation learning.
When it comes to learning features of graph topology, as illustrated in Figure 3e - Figure 3h, both RSGT-S and RSGT-P show significant performance deterioration across four all experimental datasets. Remarkably, the performance decline of RSGT-P is more severe than RSGT-S. This finding suggests that the shortest path, along with varying edge weights and types along the path, can more effectively clarify the topological dependency relationship between nodes. As anticipated, RSGT-SP performed the worst among all ablated models due to its neglect of graph topology information, depending solely on the original self-attention mechanism to learn global semantic relevance between nodes, which fails to capture the topological dependencies between nodes. This further confirms the efficacy of our structure-reinforced attention mechanism that integrates graph topological information into the Transformer architecture.
@@
IV-E Performance Impact of Parameters
Here we conduct an in-depth performance analysis with a focus on key hyperparameters.
Shortest Path Distance
Given computational resource constraints, our experimental design only incorporates the subgraph within the $spd$ hops to learn the representation of the target node. Figure 4a and Figure 4b reveal a positive correlation between the growth of $spd$ and the enhancement of model performance. This observation suggests that our proposed structure-reinforced graph transformer can effectively learn more comprehensive and informative node representations from a broader context. For the comparative experiments, we set the $spd$ to 5 for the twi-Tennis, CollegeMsg, and cit-HepTh datasets, and 2 for the sx-MathOF dataset.
Window Size
The term ’window size’ in this context refers to the quantity of consecutive snapshots preceding the target time slice that are used for model training. A small $window\ size$ could possibly miss crucial historical dynamics, while a larger size might incorporate obsolete information. As depicted in Figure 4c and Figure 4d, the model performance displays an increasing trend with the window size, suggesting that our proposed recurrent learning framework is capable of effectively employing long-term temporal data and autonomously extracting valuable information. For all baselines, the $window\ size$ is set to 5 in our experiments.
Number of Encoding Layer
The parameter $N_{layer}$, denoting the number of encoding layers, determines the depth of the SGT and impacts its ability to learn deep latent features. As shown in Figure 4e and Figure 4f, there is a direct relationship between model performance and the number of encoding layers ($N_{layer}$). Although a larger $N_{layer}$ can enhance model performance, it also increases computational costs and the risk of overfitting. Therefore, in our experimental contexts, we have set $N_{layer}$ to 4, striking a balance between computational efficiency and model performance.
Number of Attention Head
As outlined in [28], multi-head attention allows the model to attend to nodes from diverse representation subspaces at varying distances simultaneously. This is evidenced in Figure 4g and Figure 4h, where the model’s performance increases with the number of attention heads ($N_{head}$). However, an overly large number of attention heads can result in a decrease in performance. Therefore, we have chosen to set $N_{head}$ to 4.
V Related Work
@@
V-A Discrete Dynamic Graph Representation Learning
@@
Typically, structural features are initially extracted from each time slice’s snapshot using GNNs. Subsequently, a sequence model captures the evolutionary dynamics between these snapshots and updates the features accordingly. For instance, Graph Convolutional Recurrent Network (GCRN) [38] utilizes a Graph Convolutional Network (GCN) [16] to obtain node embeddings, which are then processed by a Long Short-Term Memory (LSTM) network to learn the evolutionary dynamics. Variational Recurrent Graph Neural Network (VRGNN) [39] extends GCRN by incorporating Variational Graph Auto-Encoder (VGAE) [41] to enhance the expressive power of GCRN and to capture the uncertainty in node latent representation. DySAT [40] follows a similar approach to GCRN, but calculates node representations through a joint self-attention mechanism along the structural neighborhood and temporal dynamics dimensions. Despite the intuitive nature and potential of these approaches for dynamic graph representation learning, they suffer from a lack of inductive learning capability, which is the inability to leverage the temporal information of neighboring nodes for newly introduced nodes.
@@
To address this limitation, several studies have explored methods to simultaneously learn both graph structural and dynamic features, thereby improving inductive learning. EvolveGCN [7], for example, proposes two architectures that capture the dynamism of the graph sequence by using a Recurrent Neural Network (RNN) to evolve the GCN parameters instead of the node embeddings. As the number of nodes in a dynamic graph grows across time slices, DynGEM [18] employs a deep autoencoder to progressively construct the representations of a particular snapshot from the embedding of the preceding snapshot. CoEvoSAGE [13] leverages GraphSAGE to learn the structural dependencies for each snapshot and introduces a temporal self-attention architecture to model long-range dependencies. ROLAND [8] is to easily adapt any static GNN for use with dynamic graphs. It treats the node embeddings at different GNN layers as hierarchical node states and updates them recurrently over time. DynGNN [42] integrates an RNN with a GNN for effectively capturing the evolving patterns of graphs by considering both temporal dynamics and graph topology.
V-B Continuous Dynamic Graph Representation Learning
@@
The continuous dynamic graph models edges at different temporal instances as timestamped events. These can be represented as a set of triplets $\{<v_{i},v_{j},t>\}$, where $<v_{i},v_{j},t>$ signifies the presence of an edge between nodes $v_{i}$ and $v_{j}$ at the timestamp $t$. Various representation learning methods for continuous dynamic graphs [9, 34, 35, 10] seek to understand and elucidate the evolving patterns of these graphs from such temporal event sequences.
CTDNE [9] extends random walk-based embedding approaches to cater to continuous dynamic graphs. TGAT [34], on the other hand, introduces the temporal graph attention layer to efficiently aggregate features from temporal-topological neighborhoods and captures the interactions between time and features. By stacking the temporal graph attention layers, the network treats node embeddings as time-dependent functions and can infer embeddings for both new and observed nodes as the graph evolves adaptively. CAW [35] offers Causal Anonymous Walks that are derived from temporal random walks and act as automatic retrieval mechanisms for temporal graphs, thereby bypassing the need for time-consuming selection and counting of motifs. TREND [10] leverages Hawkes processes [36] to infuse the evolutionary characteristics of temporal edges into node representations derived from continuous dynamic graphs. GraSSP [43] proposes a novel stochastic process to model link durations and their absences in continuous-time graphs. FuzzyDGL [44] leverages fuzzy logic to address uncertainties in dynamic graphs, enhancing representation quality by capturing structural and temporal correlations with improved handling of feature fuzziness.
While existing approaches have made substantial advancements, they primarily learn the evolutionary patterns of the dynamic graph in an implicit manner, without considering the various temporal states of the edges and their impact on the representation learning process. Furthermore, these methods predominantly use GNNs as their backbone to extract structural features from the graph. However, this approach suffers from an inherent over-smoothing issue, which consequently restricts the performance of the model. To address these problems, we propose a novel model, referred to as RSGT. RSGT significantly enhances the performance by targeting two key areas: the strategy for modeling the dynamic graph and the model for extracting temporal node features. Through these targeted improvements, RSGT is able to effectively bolster the overall performance.
VI Conclusion

### v3

@@ 4.3.1 Prediction task and parameter settings
Taking into consideration both the size and complexity of the datasets, we designate the dimension of node representation $d$ as 32 for the twi-Tennis, CollegeMsg, and sx-MathOF datasets, and 16 for the cit-HepTh dataset. During the testing phase, we initially generate temporal node representations using the trained model. These representations are subsequently input into a logistic regression classifier to predict the likelihood of a candidate edge being either positive or negative, assessed one snapshot at a time. For each test snapshot in $\mathcal{D}^{te}$, we adopt a 1:1 ratio of negative sampling to construct the complete test edge set. Here, 80% of the total test edges are utilized to train the classifier, while the remaining 20% are reserved for testing purposes. The scale factor $\alpha$ and the power factor $\beta$ are kept constant at 1 in all datasets. We conduct the experiment five times and report the mean classification accuracy and F1 scores, along with their respective standard deviations.
4.3.2 Comparison with competing baselines
@@
The experimental results for dynamic link prediction, as outlined in Table 2 and Table 3, reveal a comprehensive comparison amongst the various baseline methods. The sub-optimal performance is indicated by darker shading, while the most striking results are highlighted in bold. Our proposed method, RSGT, demonstrates consistently superior performance across all four experimental datasets when compared to both traditional and state-of-the-art techniques. This superiority is evident from the relative increments in Accuracy, which range from 6.10% to 15.42%, and F1 scores, showing improvements from 5.13% to 14.55%.
In comparing static and dynamic approaches, we observe that dynamic graph methods generally outperform those designed for static graphs, underscoring the importance of integrating temporal information for comprehending graph evolution. For instance, on the twi-Tennis dataset, the dynamic method CAW achieves an accuracy of 71.35% and an F1 score of 71.57%, significantly outperforming the static method GraphSAGE, which only reaches an accuracy of 62.76% and an F1 score of 62.26%. This trend is also evident in the CollegeMsg dataset, where CAW (accuracy: 73.04%, F1: 72.95%) substantially surpasses GraphSAGE (accuracy: 58.91%, F1: 60.45%). However, this pattern is not universal across all datasets. On the cit-HepTh dataset, the static method GraphSAGE (accuracy: 70.72%, F1: 71.27%) outperforms some dynamic approaches such as EvolveGCN (accuracy: 61.57%, F1: 62.42%) and CTDNE (accuracy: 49.42%, F1: 44.23%). This variability highlights the complexity of dynamic graph learning and the need for methods that can adapt to different types of temporal evolution.
The performance of different models also varies significantly across datasets, reflecting the impact of dataset characteristics on model efficacy. Social network datasets (twi-Tennis and CollegeMsg) favor models that excel in capturing short-term dependencies and frequent interactions. For example, on the twi-Tennis dataset, recent dynamic graph methods like DyGFormer show strong performance (accuracy: 75.89%, F1: 76.43%), likely due to their ability to model rapid changes in network structure. In contrast, the citation network (cit-HepTh) presents unique challenges due to its long-term dependencies and relatively slower evolution. Here, methods that can effectively capture long-range temporal dependencies, such as ConTIG, perform well (accuracy: 82.19%, F1: 82.92%). The Q&A platform dataset (sx-MathOF), with its complex interaction patterns, favors models that can balance both structural and temporal aspects of the graph. On this dataset, R-GraphSAGE demonstrates strong performance (accuracy: 81.63%, F1: 81.79%), underscoring the importance of versatile approaches that can adapt to diverse interaction patterns.
Despite the varying performance of different methods across datasets, RSGT consistently outperforms all baselines. On the twi-Tennis dataset, RSGT achieves an accuracy of 87.59% and an F1 score of 87.55%, surpassing the best baseline (DyGFormer) by 15.42% in accuracy and 14.55% in F1 score. Similar performance gains are observed across all datasets, with RSGT showing improvements of 13.74% in accuracy and 12.86% in F1 score on CollegeMsg, 6.10% in accuracy and 5.13% in F1 score on cit-HepTh, and 7.69% in accuracy and 7.47% in F1 score on sx-MathOF, compared to the best performing baselines on each dataset.
RSGT’s superior performance across all datasets can be attributed to several key innovations. Its edge temporal state modeling through a weighted multi-relation dynamic graph provides crucial heuristic information for node representation learning, while its adaptive learning mechanism mitigates the influence of erratic neighbors during recurrent updates. The structure-reinforced graph transformer architecture captures both global semantic correlations and local topological features, leading to high-quality node representations. Moreover, RSGT’s comprehensive temporal modeling, combining recurrent structure with edge temporal state modeling, effectively captures both short-term and long-term dependencies. These features collectively contribute to RSGT’s robustness, scalability, and superior inductive learning capability, enabling it to consistently outperform specialized methods across diverse dynamic graph scenarios. The model’s ability to adapt to different graph evolution patterns, coupled with its effective modeling of both structural and temporal aspects, establishes RSGT as a powerful and versatile approach for dynamic graph representation learning.
(a) Accuracy
(b) Recall
(c) Precision
(d) F1
(e) Accuracy
(f) Recall
(g) Precision
(h) F1
Figure 3: Performance of different ablated variations of RSGT. (a)-(d) show the impact of edge temporal state modeling, while (e)-(h) illustrate the effect of graph topology learning across different performance metrics and datasets.
4.4 Ablation Study
To evaluate the individual contributions of each component within the RSGT, we performed a comprehensive ablation study on dynamic link prediction tasks across four datasets. Our ablation variants can be categorized into two main groups: those related to edge temporal state modeling and those related to graph topology learning.
4.4.1 Edge Temporal State Modeling
We first examine the impact of edge temporal state modeling through the following variants:
  • 1.
    RSGT-T: Retains only edge weights without considering edge types. This variant helps us understand the importance of modeling short-term dynamics through edge type changes.
  • 2.
    RSGT-W: Retains only edge types without explicit edge weights modeling. This allows us to assess the impact of capturing long-term dynamics through the evolution of edge weights over time.
  • 3.
@@ 4.4.1 Edge Temporal State Modeling
    RSGT-TW: Uses the original dynamic graph directly without explicit modeling of edge temporal states or weights. This variant serves as a baseline to evaluate the overall impact of our temporal modeling approach.
Figures 3a-3d illustrate the performance of these variants across all datasets. We observe that RSGT-T and RSGT-W perform similarly on the twi-Tennis, ColledgeMsg, and sx-MathOF datasets, with only minor differences in their performance metrics. This suggests that both edge types and weights contribute substantially to capturing temporal state information. The full RSGT model consistently outperforms these variants across all datasets, highlighting the importance of combining both edge types and weights.
Interestingly, on the cit-HepTh dataset, RSGT and RSGT-T demonstrate clear performance superiority over RSGT-W and RSGT-TW. This can be attributed to the unique characteristics of the cit-HepTh dataset, which is a citation network. In such networks, edges (citations) are inherently stable once formed, rarely disappearing over time. This stability emphasizes the importance of long-term dynamics in the graph structure.
4.4.2 Graph Topology Learning
To assess the importance of graph topology learning, we evaluate the following variants:
  • 1.
@@ 4.4.2 Graph Topology Learning
When it comes to learning features of graph topology, as illustrated in Figure 3e-3h, both RSGT-S and RSGT-P show significant performance deterioration across four all experimental datasets. Remarkably, the performance decline of RSGT-P is more severe than RSGT-S. This finding suggests that the shortest path, along with varying edge weights and types along the path, can more effectively clarify the topological dependency relationship between nodes. As anticipated, RSGT-SP performed the worst among all ablated models due to its neglect of graph topology information, depending solely on the original self-attention mechanism to learn global semantic relevance between nodes, which fails to capture the topological dependencies between nodes. This further confirms the efficacy of our structure-reinforced attention mechanism that integrates graph topological information into the Transformer architecture.
@@
Figure 4: Impact of key hyperparameters on RSGT performance across four datasets (twi-Tennis, CollegeMsg, cit-HepTh, sx-MathOF). The subplots show the effect of (a-b) shortest path distance ($spd$), (c-d) window size, (e-f) number of encoding layers ($N_{layer}$), and (g-h) number of attention heads ($N_{head}$) on model performance. The x-axis in each subplot represents the varying values of the respective parameter, and the y-axis shows the corresponding performance metric in percentage.
4.5 Performance Impact of Parameters
To optimize the performance of our proposed RSGT model, we conducted a comprehensive analysis of four key hyperparameters: shortest path distance ($spd$), window size, number of encoding layers ($N_{layer}$), and number of attention heads ($N_{head}$). This analysis aims to understand the impact of each parameter on model performance and determine their optimal values for our experiments.
Shortest Path Distance The $spd$ parameter determines the subgraph size used to learn the representation of the target node. We varied $spd$ from 1 to 9 and evaluated its impact on model performance. As shown in Figures 4a and 4b, there is a positive correlation between $spd$ and model performance across all datasets. This trend suggests that our structure-reinforced graph transformer effectively leverages information from a broader context to learn more comprehensive and informative node representations. The performance improvement is particularly notable for the sx-MathOF dataset, which may be due to its more complex network structure requiring a larger neighborhood for effective representation learning. However, we observe that the improvement tends to plateau after $spd=5$ for most datasets, likely due to the inclusion of less relevant information from distant nodes. Based on these results and considering computational efficiency, we set $spd$ to 5 for twi-Tennis, CollegeMsg, and cit-HepTh datasets, and 2 for the sx-MathOF dataset in our comparative experiments.
Window Size The window size parameter defines the number of consecutive snapshots preceding the target time slice used for model training. We experimented with window sizes ranging from 1 to 9. Figures 4c and 4d demonstrate that model performance generally improves as the window size increases. This trend indicates that our recurrent learning framework effectively utilizes long-term temporal data and autonomously extracts valuable historical information. The improvement is most pronounced for the cit-HepTh dataset, which aligns with the nature of citation networks where long-term dependencies are crucial. Conversely, the CollegeMsg dataset shows less sensitivity to increased window size, possibly due to the more transient nature of messaging interactions. To balance performance gains with computational costs, we set the window size to 5 for all datasets in our experiments.
Number of Encoding Layer The $N_{layer}$ parameter determines the depth of the structure-reinforced graph transformer and impacts its ability to learn deep latent features. We varied $N_{layer}$ from 1 to 9 in our experiments. Figures 4e and 4f reveal a general trend of improved performance with increased $N_{layer}$. This suggests that deeper models can capture more complex patterns in the dynamic graphs. However, we observe diminishing returns and even slight performance degradation for some datasets (e.g., cit-HepTh) at higher $N_{layer}$ values. This could be attributed to overfitting or the vanishing gradient problem in very deep networks. Considering the trade-off between model performance and computational complexity, we chose $N_{layer}=4$ for our experiments, which provides a good balance across all datasets.
Number of Attention Head Multi-head attention allows the model to focus on different representation subspaces simultaneously. We experimented with $N_{head}$ values of 1, 2, 4, 8, and 16. As shown in Figures 4g and 4h, increasing $N_{head}$ generally improves model performance, particularly from 1 to 4 heads. This aligns with the findings of [30], demonstrating the benefit of attending to multiple representation subspaces. However, we observe that performance tends to plateau or even slightly decrease beyond 4 heads for most datasets. This could be due to increased model complexity leading to overfitting, especially on smaller datasets. Based on these results, we set $N_{head}=4$ in our experiments, which provides optimal performance across datasets while maintaining computational efficiency.
5 Related Work
@@
5.1 Discrete Dynamic Graph Representation Learning
@@
Typically, structural features are initially extracted from each time slice’s snapshot using GNNs. Subsequently, a sequence model captures the evolutionary dynamics between these snapshots and updates the features accordingly. For instance, Graph Convolutional Recurrent Network (GCRN) [43] utilizes a Graph Convolutional Network (GCN) [39] to obtain node embeddings, which are then processed by a Long Short-Term Memory (LSTM) network to learn the evolutionary dynamics. Variational Recurrent Graph Neural Network (VRGNN) [44] extends GCRN by incorporating Variational Graph Auto-Encoder (VGAE) [46] to enhance the expressive power of GCRN and to capture the uncertainty in node latent representation. DySAT [45] follows a similar approach to GCRN, but calculates node representations through a joint self-attention mechanism along the structural neighborhood and temporal dynamics dimensions.
@@
Despite the intuitive nature and potential of these approaches for dynamic graph representation learning, they suffer from a lack of inductive learning capability, which is the inability to leverage the temporal information of neighboring nodes for newly introduced nodes. To address this limitation, several studies have explored methods to simultaneously learn both graph structural and dynamic features, thereby improving inductive learning. EvolveGCN [7], for example, proposes two architectures that capture the dynamism of the graph sequence by using a Recurrent Neural Network (RNN) to evolve the GCN parameters instead of the node embeddings. As the number of nodes in a dynamic graph grows across time slices, DynGEM [19] employs a deep autoencoder to progressively construct the representations of a particular snapshot from the embedding of the preceding snapshot. CoEvoSAGE [13] leverages GraphSAGE to learn the structural dependencies for each snapshot and introduces a temporal self-attention architecture to model long-range dependencies. ROLAND [8] is designed to easily adapt any static GNN for use with dynamic graphs. It treats the node embeddings at different GNN layers as hierarchical node states and updates them recurrently over time. DynGNN [47] integrates an RNN with a GNN for effectively capturing the evolving patterns of graphs by considering both temporal dynamics and graph topology.
5.2 Continuous Dynamic Graph Representation Learning
@@
The continuous dynamic graph models edges at different temporal instances as timestamped events. These can be represented as a set of triplets $\{\langle v_{i},v_{j},t\rangle\}$, where $\langle v_{i},v_{j},t\rangle$ signifies the presence of an edge between nodes $v_{i}$ and $v_{j}$ at the timestamp $t$. Various representation learning methods for continuous dynamic graphs [9, 15, 16, 10] seek to understand and elucidate the evolving patterns of these graphs from such temporal event sequences.
Continuous dynamic graph representation learning methods can be broadly categorized into two main approaches: Sequence-based modeling methods, and Advanced temporal modeling methods. Sequence-based modeling methods aim to capture temporal-topological information by treating the graph evolution as a sequence of events. CTDNE [9] generalizes random walk-based embedding techniques to continuous dynamic graphs, effectively capturing local temporal patterns but potentially struggling with long-range dependencies. Attention-based methods like TGAT [15] and CAW [16] leverage self-attention mechanisms to capture temporal-topological interactions. TGAT introduces a temporal graph attention layer to aggregate features from temporal-topological neighborhoods, while CAW proposes Causal Anonymous Walks as an automatic retrieval mechanism for temporal graphs. These methods excel at capturing complex temporal dependencies but may face computational challenges with very large graphs. DyGFormer [41] employs a Transformer-based architecture with neighbor co-occurrence encoding and patching techniques to learn from nodes’ historical first-hop interactions, effectively capturing both short-term and long-term dependencies.
Advanced temporal modeling methods employ sophisticated techniques to model the continuous evolution of graphs. TREND [10] leverages Hawkes processes to incorporate the evolutionary characteristics of temporal edges into node representations. R-GSAGE [17] uses neural ordinary differential equations to model the continuous dynamic evolution of node embedding trajectories, allowing for a more nuanced representation of temporal dynamics. GraSSP [18] proposes a novel stochastic process to model link durations and their absences in continuous-time graphs. FuzzyDGL [48] applies fuzzy logic to address uncertainties in dynamic graphs, enhancing representation quality by capturing structural and temporal correlations with improved feature fuzziness handling. ConTIG [20] integrates a recurrent structure into GraphSAGE to jointly explore structural and temporal patterns while maintaining a lightweight architecture, making it particularly efficient for large-scale dynamic graphs.
Despite these advancements, existing approaches still face several challenges. Many methods learn evolutionary patterns implicitly, without explicitly considering the various temporal states of edges and their impact on node representations. Furthermore, the prevalent use of GNNs as the backbone for structural feature extraction often leads to over-smoothing issues, limiting model performance. Additionally, most current approaches struggle to effectively balance the modeling of both short-term and long-term temporal dependencies, especially in graphs with diverse temporal scales. Our proposed RSGT addresses these limitations by introducing a novel edge temporal state modeling technique and a structure-reinforced graph transformer. It allows for explicit modeling of edge dynamics while effectively capturing both local and global structural information, thereby overcoming the challenges faced by existing methods.
6 Conclusion and Future Work